What is 11001100? (worked example)

1

1001

100

Answer: - 1.100 (base 2) * 2 ^ 2 = - 1.5 * 4 = 6

  1. What is 11010010?

               How 2.5 will be represented? (worked example)

0

1000

010

2.5 = 10.1 (base 2) = 1.01 (base 2) * 2^1

Answer: 01000010

  1. How 10.0 will be represented?
  1. How -0.25 will be represented?

Answer:
Answer #1

The first bit is sign bit 0 means its positive 1 means the number is negative

Next 4 bit represents the integer part

The last 3 bit represents the decimal part

==========================================================

a) What is 11010010?
1 = Sign bit

1010 = 8 + 2 = 10
010 = 0*2^-1 + 1*2^-2 + 0*2^-3 = 0 + 0.25 + 0

So the final answer is -10.25

==========================================================

b) How 10.0 will be represented?

sign part = 0
integer part = 1010
decimal part = 000

So 10.0 will be represented as 01010000

==========================================================
c) How -0.25 will be represented?

sign part = 1
integer part = 0000
decimal part = 010

So -0.25 will be represented as 10000010